OpenStudy (anonymous):

A gallon of paint costs $12.99 and covers 400 square feet. You are painting two coats on the walls and ceiling of a rectangular room that is 30 feet long, 18 wide, and 8 feet high. The floor is not painted.

Directrix (directrix):

I guess we need to get the area of the 4 walls of the room and then the area of the ceiling. So, it's like the total area of a rectangular solid minus the area of one base (floor not painted).

OpenStudy (anonymous):

It says two coats. Does that have to do with anything?

Directrix (directrix):

Yes, but let's get the area to be painted first.

Directrix (directrix):

Here's the room.

Directrix (directrix):

Area to be painted = 2* [18 * 8] + 2 * [30 * 8] + 1* [30 * 18]. Before we add these, see if they make sense.

OpenStudy (anonymous):

Alright. Hold on.

OpenStudy (anonymous):

Okay, I follow

OpenStudy (anonymous):

Here are the questions needed to answer: a. How many gallons are needed? b. How much will it cost?

Directrix (directrix):

2* [18 * 8] + 2 * [30 * 8] + 1* [30 * 18] = 1308 square feet of area. Each square foot of area is to be painted twice. So, there is the equivalent of 2616 square feet of area for which to buy paint.

OpenStudy (anonymous):

7?

Directrix (directrix):

One gallon is to 400 square feet as x gallons is to 2616 1/400 = x/2616 x = 6.54 which is 7 gallons of paint

Directrix (directrix):

And, the cost is ?

OpenStudy (anonymous):

90.93

Directrix (directrix):

Yes.

Directrix (directrix):

Well, I guess that's about it for now, yes?

OpenStudy (anonymous):

Yes, thank you :) @Directrix